OpenStudy (anonymous):

A jet boat is used to take canoers 200 miles upriver to their starting point. As the driver of the jet boat, Jeremy knows that it will take him 5 hours to get the canoers upstream, but only 4 hours to get the boat back downstream after dropping the canoers off, as long as he keeps the boat moving at a constant speed. How fast would the boat be moving in still water? A. 5 mph B. 40 mph C. 45 mph D. 50 mph

OpenStudy (nurali):

200/5 = 40 mph 200/4 = 50 mph Let x be the speed in still waters and y the speed of current. x + y = 50 x - y = 40 ADD 2x = 90 x = 45 mph

OpenStudy (anonymous):

thank you so much , makes much more sense c: @Nurali

OpenStudy (nurali):

My Pleasure.

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

Rhnak you so much (: @Nurali

OpenStudy (anonymous):

thank*